# The Renormalization Group for Large-Scale Structure: Origin of Galaxy Stochasticity

### v1

@@ 5 Conclusion
We have used the general effective action for large-scale structure, Eq. (2.2), to derive the “RG-LSS” equations which describe the running of bias and stochastic contributions to the clustering of LSS tracers under a change of the renormalization scale. This extends the analysis of Paper 1 [10], which focused on bias, i.e. terms linear in the current $J$, to arbitrary powers of $J$. In Sec. 3, we make use of the Wilson-Polchinski formalism to derive how those terms are sourced by bias operators and lower-order-in-$J$ stochastic parameters. We provide a very general master formula Eq. (3.36) for the evolution of a general stochastic parameter as a function of the cutoff $\Lambda$, which simplifies to Eq. (3.40) for the shot-noise contributions to $n$-point functions. We also analyse in Sec. 4 solutions to the RG-flow including stochastic parameters.
In this work we continue to follow the philosophy introduced in Paper 1, which keeps the smoothing cutoff (i.e., the renormalization scale) finite, instead of subtracting the leading order contributions by taking the large-scale limit $\Lambda\to 0$ as considered in usual EFT of LSS analyses. The finite-$\Lambda$ allows for the derivation of the RG equations using the Wilson-Polchinski framework, which may account for a yet-to-be-determined amount of extra information from the resummation of part of higher-loop contributions as in the usual context of quantum field theory. We find that, different than for the bias parameters, the RG group equations for the stochastic parameters are nonlinear, which may suggest that some relevant information may be absorbed by the their RG flow. Furthermore, non-trivial critical points may arise from the RG flow structure of the bias and stochastic parameters. These are directions we aim to investigate in the future. Another interesting generalization to be understood in light of the RG equations is the presence of primordial non-Gaussianities. Non-Gaussianities lead to new types of vertices for which the RG equations (and their inbuilt resummation) may be relevant.
@@ Acknowledgments
HR is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2094 ‘ORIGINS’. (No. 390783311). We thank Giovanni Cabass, Mathias Garny, Charalampos Nikolis, and Marko Simonovic for discussions and Charalampos Nikolis and Beatriz Tucci for feedback on the paper.
@@ A.3 The $\mathcal{S}_{OO^{\prime}}$ integrals
Notice that $\sigma^{2}_{\bm{p},-\bm{k}-\bm{p}}$ will also source higher-derivative terms. Terms proportional to $P_{\rm shell}(k)$ will again be zero when integrated with $J_{\Lambda}(\bm{k})$ due to the orthogonality condition Eq. (3.3). The only non-zero contribution at this level is then $\left(\mathcal{S}^{11}_{\delta^{2}\delta^{2}}\right)_{(2)_{\rm leg}}$, but it does not contribute by the arguments given in App. B.
@@ Appendix B On the general structure of the shell contributions at the $n$-point function level
This appendix is dedicated to understand the general structure of the sourcing terms $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}$ at different orders in $J$. For that, we take the approach of calculating the corrections to $n$-point functions  [^10] An analysis at the level of the partition function leads to intricate momenta structure of the shell terms, such that the $n$-point functions is the easiest way to understand it.. The main conclusions of this appendix can be summarized as: The terms of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{\dots 1\dots}$ are zero when calculating $n$-point functions due to their kinematic structure. Thus, the leading contribution comes from the terms of the type $\mathcal{S}_{O_{1}O_{2}\dots Op}^{22\dots 2}$.
@@ B.1 The contribution of $\mathcal{S}_{OO^{\prime}}$ to $J^{2}$ terms
Comparing with Eq. (B.1), evaluated at leading order $\delta[\delta^{(1)}_{\Lambda}]\to\delta^{(1)}_{\Lambda}$, we see that Eq. (B.7) has a similar structure to the $P_{\epsilon,\delta^{2}}$ contribution, but with $\delta_{\rm D}(\bm{x}-\bm{x}^{\prime})$ replaced with $\xi_{\rm shell}(\bm{x}-\bm{x}^{\prime})$. We obtain the contribution to the 4-point function which corresponds to Eq. (B.3):  [^11] Notice we can use the absence of preferred directions in the initial conditions, and the assumption of a spherical filter $W_{\Lambda}$, to write $\xi_{\rm shell}(\bm{x}-\bm{x}^{\prime})=\xi_{\rm shell}(|\bm{x}-\bm{x}^{\prime}|)$ $=\int_{\bm{k}\in\rm shell}P_{\rm L}(k)\exp[\operatorname{i}\bm{k}\cdot(\bm{x}-\bm{x}^{\prime})]$ $=\frac{1}{2\pi^{2}}\Lambda^{2}\lambda P_{\rm L}(\Lambda)\frac{\sin(\Lambda r)}{\Lambda r}\Big{|}_{r=|\bm{x}-\bm{x}^{\prime}|}=\sigma_{\rm shell}^{2}\frac{\sin(\Lambda r)}{\Lambda r}\Big{|}_{r=|\bm{x}-\bm{x}^{\prime}|}\,.$ (B.8) In the second line, we have additionally assumed a shell of infinitesimal width $\lambda$. This shows the expected structure: $\xi_{\rm shell}(0)=\sigma_{\rm shell}^{2}$, and it decays on the scale $r\sim 1/\Lambda$.
@@
In order to connect Eq. (B.3) and Eq. (B.9), we now assume that the 4-point function is integrated against a test function $f(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3},\bm{x}_{4})$ which defines the configurations or quadrilateral bins for which the 4-point function is measured. Given the assumption about the support of the current $J_{\Lambda}(\bm{k})$, we can assume that, for all of its arguments, $f$ only has support on large scales, $k\lesssim k_{\rm max}\ll\Lambda$. Eq. (B.3) straightforwardly yields
@@ B.2 The structure of the $n$-point functions for the most general case
One can use a very similar argument to the one presented in App. B.1 to show that any shell contribution of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{\dots 1\dots}$ is zero. Considering whatever $n$-point function contains this term at tree level, similarly to Eq. (B.11) and Eq. (B.14), this term will contain the structure

### v2

@@ 5 Conclusion
We have used the general effective action for large-scale structure, Eq. (2.2), to derive the “RG-LSS” equations which describe the running of bias and stochastic contributions to the clustering of LSS tracers under a change of the renormalization scale. This extends the analysis of Paper 1 [10], which focused on bias, i.e. terms linear in the current $J$, to arbitrary powers of $J$. Our result therefore is the most general RG running for the leading-in-derivative terms of the EFT of LSS. In Sec. 3, we make use of the Wilson-Polchinski formalism to derive how those terms are sourced by bias operators and lower-order-in-$J$ stochastic parameters. We provide a general master formula Eq. (3.36) for the evolution of any stochastic or bias parameter as a function of the cutoff $\Lambda$, which simplifies to Eq. (3.40) for the shot-noise contributions to $n$-point functions. We also analyse in Sec. 4 solutions for the RG-flow including stochastic parameters.
In this work we continue to follow the philosophy introduced in Paper 1, which keeps the smoothing cutoff (i.e., the renormalization scale) finite, instead of subtracting the leading order contributions by taking the large-scale limit $\Lambda\to 0$ as considered in usual EFT of LSS analyses (see Sections 3 and 4 of Paper 1 for a comparison between both finite-$\Lambda$ and $n$-point-function renormalization schemes). [^9] For the renormalization of stochastic terms, see e.g. Appendix D of [30, 31]. The finite-$\Lambda$ approach allows for the derivation of the RG equations using the Wilson-Polchinski framework, which may account for a yet-to-be-determined amount of extra information from the resummation of part of higher-loop contributions as in the usual context of quantum field theory, where large logarithms are resummed. We find that, different than for the bias parameters, the RG equations for the stochastic parameters are nonlinear, which may suggest that some relevant information may be absorbed by their RG flow. The discussion of whether more (non-perturbative) information can be extracted from the RG equations for the LSS coefficients is deferred for a future publication. Furthermore, non-trivial critical points may arise from the RG flow structure of the bias and stochastic parameters. These are directions we aim to investigate in the future. Another interesting generalization to be understood in light of the RG equations is the presence of primordial non-Gaussianities [32]. Non-Gaussianities lead to new types of vertices for which the RG equations (and their inbuilt resummation) may be relevant.
@@ Acknowledgments
HR is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC 2094 ‘ORIGINS’. (No. 390783311). We thank Giovanni Cabass, Mathias Garny, Matt Lewandowski, Mehrdad Mirbabayi, Charalampos Nikolis, and Marko Simonovic for discussions and Charalampos Nikolis and Beatriz Tucci for feedback on the paper.
@@ A.3 The $\mathcal{S}_{OO^{\prime}}$ integrals
Notice that $\sigma^{2}_{\bm{p},-\bm{k}-\bm{p}}$ will also source higher-derivative terms. Terms proportional to $P_{\rm shell}(k)$ will again be zero when integrated with $J_{\Lambda}(\bm{k})$ due to the orthogonality condition Eq. (3.3). The only non-zero contribution at this level is then $\left(\mathcal{S}^{11}_{\delta^{2}\delta^{2}}\right)_{(2)_{\rm leg}}$, but it does not contribute in the limit of soft external momenta by the arguments given in App. B.
@@ Appendix B On the general structure of the shell contributions at the $n$-point function level
This appendix is dedicated to understand the general structure of the sourcing terms $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}$ at different orders in $J$. For that, we take the approach of calculating the corrections to $n$-point functions  [^10] An analysis at the level of the partition function leads to intricate momenta structure of the shell terms, such that the $n$-point functions is the easiest way to understand it.. The main conclusions of this appendix can be summarized as: The terms of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{\dots 1\dots}$ lead to nonlocal interactions in the partition function. However, in the limit when external momenta are much smaller than the scale $\Lambda$, these terms vanish due to their kinematic structure. In this limit, the leading contribution comes from the terms of the type $\mathcal{S}_{O_{1}O_{2}\dots Op}^{22\dots 2}$.
@@ B.1 The contribution of $\mathcal{S}_{OO^{\prime}}$ to $J^{2}$ terms
Comparing with Eq. (B.1), evaluated at leading order $\delta[\delta^{(1)}_{\Lambda}]\to\delta^{(1)}_{\Lambda}$, we see that Eq. (B.7) has a similar structure to the $P_{\epsilon,\delta^{2}}$ contribution, but with $\delta_{\rm D}(\bm{x}-\bm{x}^{\prime})$ replaced with $\xi_{\rm shell}(\bm{x}-\bm{x}^{\prime})$. That is, Eq. (B.7) is nonlocal, while the interaction written in the partition function is local. We obtain the contribution to the 4-point function which corresponds to Eq. (B.3):  [^11] Notice we can use the absence of preferred directions in the initial conditions, and the assumption of a spherical filter $W_{\Lambda}$, to write $\xi_{\rm shell}(\bm{x}-\bm{x}^{\prime})=\xi_{\rm shell}(|\bm{x}-\bm{x}^{\prime}|)$ $=\int_{\bm{k}\in\rm shell}P_{\rm L}(k)\exp[\operatorname{i}\bm{k}\cdot(\bm{x}-\bm{x}^{\prime})]$ $=\frac{1}{2\pi^{2}}\Lambda^{2}\lambda P_{\rm L}(\Lambda)\frac{\sin(\Lambda r)}{\Lambda r}\Big{|}_{r=|\bm{x}-\bm{x}^{\prime}|}=\sigma_{\rm shell}^{2}\frac{\sin(\Lambda r)}{\Lambda r}\Big{|}_{r=|\bm{x}-\bm{x}^{\prime}|}\,.$ (B.8) In the second line, we have additionally assumed a shell of infinitesimal width $\lambda$. This shows the expected structure: $\xi_{\rm shell}(0)=\sigma_{\rm shell}^{2}$, and it decays on the scale $r\sim 1/\Lambda$.
@@
In order to connect Eq. (B.3) and Eq. (B.9), we now assume that the 4-point function is integrated against a test function $f(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3},\bm{x}_{4})$ which defines the configurations or quadrilateral bins for which the 4-point function is measured. If we asssume that the current $J_{\Lambda}(\bm{k})$ only has support on large scales, $k\lesssim k_{\rm max}\ll\Lambda$, Eq. (B.3) straightforwardly yields
@@ B.2 The structure of the $n$-point functions for the most general case
One can use a very similar argument to the one presented in App. B.1 to show that any shell contribution of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{\dots 1\dots}$ is zero in the limit of soft external momenta. Considering whatever $n$-point function contains this term at tree level, similarly to Eq. (B.11) and Eq. (B.14), this term will contain the structure